An antelope is being pulled by two forces whose vector-component expressions are: F1 = 6Ni + -3N) + -1.1 Nk F2 = -3.8Ni + 7.1 Nj + 2.8Nk What is the angle in degrees) between these two vectors?



1
Expert's answer
2021-10-03T13:23:44-0400
"\cos a=\frac{-6(3.8)-3(7.1)-1.1(2.8)}{\sqrt{3.8^2+7.1^2+2.8^2}\sqrt{6^2+3^2+1.1^2}}=-0.814\\a=\arccos{(-0.814)}=144.5\degree"